Consider the following data:

X468919Y1216253039

 Which of the following statements is correct?


1
Expert's answer
2021-08-13T09:43:02-0400


By using Excel



=SLOPE(B2:B6,A2:A6)

slope = 1.73

=INTERCEPT(B2:B6,A2:A6)

intercept = 8.53

"Y=1.73x+8.53 \\

Y(10) = 1.73 \times 10 + 8.53 = 25.83 \\
R^2=0.855"

(from trend line)

Answer: B. When X is 10, then the estimated sales will be 25.83 is correct